the width of a rectangle is 3 inches less than twice its lenght. If
perimeter of the rectangle is 36 inches, find its lenght and width.

Answer by checkley77(12844)

W=L-3
2W+2L=36
2(L-3)+2L=36
2L-6+2L=36
4L=36+6
4L=42
L=42/4
L=10.5 IN.
W=10.5-3=7.5 IN.
PROOF:
2*7.5+2*10.5=36
15+21=36
36=36